# Deep learning modelling of manufacturing and build variations on multi-stage axial compressors aerodynamics

### v4

@@ Abstract
Applications of deep learning to physical simulations such as Computational Fluid Dynamics have recently experienced a surge in interest, and their viability has been demonstrated in different domains. However, due to the highly complex, turbulent and three-dimensional flows, they have not yet been proven usable for turbomachinery applications. Multi-stage axial compressors for gas turbine applications represent a remarkably challenging case, due to the high-dimensionality of the regression of the flow-field from geometrical and operational variables. This paper demonstrates the development and application of a deep learning framework for predictions of the flow field and aerodynamic performance of multi-stage axial compressors. A physics-based dimensionality reduction unlocks the potential for flow-field predictions, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. Compared to traditional ”black-box” surrogate models, it provides explainability to the predictions of overall performance by identifying the corresponding aerodynamic drivers. This is applied to model the effect of manufacturing and build variations, as the associated performance scatter is known to have a significant impact on $CO_{2}$ emissions, therefore posing a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark, in real-time, for an industrially relevant application. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance with actionable and explainable data.
@@
journal: Data-centric Engineering
@@
[SE]organization=Siemens Energy, addressline=1 Waterside South, city=Lincoln, postcode=LN5 7FD, country=United Kingdom
@@
[UL]organization=University of Lincoln, addressline=Brayford Way, city=Lincoln, postcode=LN6 7DL, country=United Kingdom
@@ 1 Introduction
The highly complex, turbulent and three-dimensional flow-field of multi-stage axial compressors has traditionally been one the main challenges for the application of CFD (Computational Fluid Dynamics) in the turbomachinery industry. In this context, the flow field consists of the aerodynamic variables of interest (such as pressure, temperature, velocities etc.) predicted on a numerical discretizations of the domain (i.e. mesh or grid), solving the Reynolds-Averaged Navier Stokes equations (RANS). Advancements in modelling and numerical methods have progressively improved its accuracy and computational cost, making CFD analyses an integral part of industrial design processes and greatly improving the design capabilities of gas turbine manufacturers. This has led to an interest in the development of robust design strategies to account for real world effects in the design and analyses, which are known to affect the overall performance as well as the operability of the compressor. Gas turbine manufacturers have gathered vast amount of operational data over the years, enabling them to establish best practices and design guidelines for their manufacturing and build processes. These guidelines, often consider a trade-off between product cost and engine performance, with tolerance ranges specified accordingly. However, the impact on performance used to define these ranges is typically based on previous experience and simplified correlations. The identification of the impact of manufacturing and build uncertainties on the operational performance can be used to both drive improvements in the manufacturing technology, and inform the design strategies to produce robust designs desensitized to these variabilities.
@@
The challenge: Characterizing the impact of all the relevant manufacturing and build variations which are expected to affect the overall performance of a given compressor design, can be achieved by engine build-specific CFD models. While having demonstrated to provide excellent agreement with test data [1], these models are not traditionally used as part of the manufacturing and build process as a day-to-day occurrence, due to the associated computational cost and requirement for specialized engineers to carry out the analyses.
@@
Our contribution: This paper demonstrates the development and application of a deep learning framework previously introduced by the authors [2] for predictions of the flow field and aerodynamic performance of multi-stage axial compressors, demonstrating the scalability from a proof of concept to a significantly more computationally expensive application of great industrial relevance. A physics-based dimensionality reduction unlocks the potential to use supervised learning approaches for flow-field predictions, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. The proposed architecture also has a competitive advantage compared to traditional ”black-box” surrogate models, as it provides explainability to the predictions of variations in overall performance by identifying the corresponding aerodynamic drivers. This is applied to model the effect of manufacturing and build variations, as the associated scatter in efficiency is known to have a significant impact on the overall performance and $CO_{2}$ emissions of the gas turbine, therefore posing a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark, in real-time, for an industrially relevant application. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance with actionable and explainable data. This unlocks the capability to model each engine build avoiding the high computational costs, long lead times, specialized personnel requirements of traditional CFD analyses, and reducing the requirement for expensive physical tests.
@@ 1.1 Effect of manufacturing and build variations on the aerodynamic performance
The impact of manufacturing and build variations on the overall performance of gas turbines is known to be significant [3] [4], with the axial compressor being a large contributor. For a given design, compressor efficiency is affected by a combination of in-tolerance variations, which can lead to a substantial increase in $CO_{2}$ emissions compared to the design target. Therefore, accurately predicting the performance of each engine build by modelling the impact of manufacturing and build variations on engine performance within a short timescale is of great industrial and environmental importance. The focus of the current work is on tip clearance and surface roughness variations, two of the main sources of performance variability, but the framework is readily generalisable to other manufacturing variations. A detailed review of the effect of the various manufacturing and build variations on the compressor aerodynamic performance, and the associated modelling strategies has been previously been provided by the authors [1]. In the following sections, a brief overview is included to provide the readers with the necessary background to interpret the results presented in this paper.
@@ 1.1.1 Tip Clearance variation background
The term tip clearance refers to the radial distance between the stationary and rotating components within the compressor, such as the rotor blades and casing, or stator blades and rotor hub. Generally, the tip clearance must be large enough to prevent rubbing during operation, but small enough to avoid excessive aerodynamic losses. Axial compressors are designed to operate at a defined tip clearance, which can vary across stages. During manufacturing, the ”cold clearance” is specified and measured as part of the build process in stationary conditions. However, during operation, the ”hot clearance” can be significantly smaller due to thermal expansion and centrifugal effects. The cold clearance values have a target value and tolerance range, set by gas turbine manufacturers, based on trade-offs between undesirable rubbing and minimizing tip leakage losses. For a given ”cold clearance”, the corresponding ”hot clearance” can be calculated using either analytical models or operational experience. The ”hot clearance” is the actual clearance found during operation, and therefore is used in the CFD and machine learning models presented in this paper. The aerodynamic mechanisms driving tip leakage flows have been extensively researched over the years [5] and can be modeled analytically with varying degrees of fidelity. Standard practice in both academia and industry, is to use steady-state single passage CFD calculations to model the effects of tip clearance variations on overall performance [6], which are used as ground-truth for this study. This assumes that circumferential variations in tip clearance can be neglected. Overall, non-axisymmetric tip clearance variations are known to affect the overall performance and surge margin [7] and are of relevance especially for in-service degradation. However, for manufacturing variations they can be neglected [2] and single passage models can be used for the analyses.
@@ 1.1.2 Surface roughness variation background
Traditionally, CFD analyses are performed assuming aerofoil and gas path surfaces to be aerodynamically smooth. This assumption is often not valid for production components, which are bound to have non-negligible roughness due to coating selection or manufacturing methods. A trade-off can be identified between production cost and roughness, as smoother finishes tend to be progressively more expensive. Therefore, it is of paramount importance to be able to characterize the impact of different surface roughness levels on the aerodynamic performance. This is particularly relevant when comparing engine test data from different builds with coating from different suppliers, as they can potentially have significantly different roughness. Experimental work has shown that increased roughness on the aerofoil surface has a detrimental effect on efficiency. One possible way to model analytically the effect of surface roughness is to consider the average $R_{a}$ for each blade row, which can be measured for a given set of blades. The $R_{a}$ can then be converted to an equivalent sand grain roughness that can be prescribed in the CFD model on the airfoils surfaces.
@@ 1.2 Machine Learning applications to turbomachinery
In recent years, there has been a significant amount of research into the utilization of machine learning techniques for turbomachinery applications. These include predicting the aerodynamic performance due to tip clearance variations [8] and design modifications [9], as well as aeromechanic predictions for both forced response [10] and flutter [11]. As often the output variables are predicted directly from the input variables with a ”black box” approach, these methodologies have varying degrees of accuracy and generalisability, depending on the complexity of the application. However, the ability to accurately predict variations in target variables for turbomachinery in a manner that is both reliable and generalisable, is contingent upon the accuracy of flow field predictions. It is only through precise flow field predictions that the overall performance figure of interest can be calculated through the application of relevant physical equations that govern the dynamics. The potential benefits of using machine learning approaches for predicting full flow fields, have been demonstrated in some simplified cases in literature for 2D aerofoils on both cartesian [12] [13] [14] and unstructured [15] grids. CNNs have been demonstrated to predict the flow field downstream of a single fan row [16], which was then used for noise predictions applications. A similar approach was used also by Rao et. al. [17] for the semantic segmentation of an aero-engine intake using U-Net and U-Net++ architectures. Graph neural networks [18] [19] have also been applied to 3D turbomachinery, with a single row compressor [20] and turbine [21], which suggested potential challenges related to scalability and practicality for industrial applications. For instance, the computational cost associated with training, and the size of the required model, make the approach impractical for multi-stage compressors. These applications typically use computational meshes in the order of magnitude of 10 million nodes. Furthermore, storing the CFD data required for training and future simulations would require several petabytes, for a typical engine manufacturer, which is not practical. Purely data-driven approaches typical of supervised learning require large amount of data, so that the network can learn the solution manifold. A recent review [22] has highlighted the potential of other approaches to mitigate these issues such as Physics-Informed Neural Networks (PINNs), Reinforcement Learning (RL), and Generative Adversarial Networks (GANs) [23]. However, none of those methods can currently provide accurate, generalisable and scalable solutions to complex 3D turbomachinery applications.
1.3 Physics-based dimensionality reduction
@@
To address these challenges, we introduce a pre-processing step to extract only the relevant engineering data from the 3D computational domain, such as: blade-to-blade planes, blade surfaces, axial cuts, loading distributions, radial distributions, stage-wise performance, and overall performance. In most cases, these provide sufficient information to the aerodynamicist for design and analyses activities, while close examination of the 3D flow field of a CFD solution is generally only necessary in non-standard cases. Once the relevant engineering data has been extracted from the computational domain, it can be interpolated onto a simplified grid with a resolution that is deemed acceptable for the application in question. The regression problem is therefore reformulated, from predicting the relevant variables on an unstructured computational grid in the order of magnitude of $\sim 10^{7}$ nodes, to a structured grid in the order of magnitude of $\sim 10^{5}$ nodes. Only selected fundamental variables are predicted, while the derived ones can be computed using the appropriate physical equations, significantly reducing the computational cost without sacrificing accuracy, as only the necessary data required to calculate the target variables is retained. Targeting the flow field predictions and using the relevant physical equations to calculate the corresponding overall performance renders the methodology generalisable, while filtering only relevant parts of the CFD solution makes the methodology scalable to complex industrial applications. This physics-based dimensionality reduction, unlocks the potential to use supervised learning approaches for turbomachinery applications, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. The authors first introduced this approach in a previous work for a single stage compressor application [2], which is now extended to a multi-stage compressor application, with a computational domain one order of magnitude larger. The framework developed in the previous paper is demonstrated to be translated from an academic test case, to a complex real-world application with significant engineering relevance. While the focus of the current work is on multi-stage axial compressors, the framework is directly applicable to any other turbomachinery application, such as axial turbines.
@@ 2.1 Data Generation
The ground-truth data used for training are the CFD results for the configuration of interest. This computational model consisted of 10 stages, with mixing plane interfaces. To reduce the computational cost of the CFD analysis, rotational periodicity is assumed, and only a single passage needs to be considered instead of the full annulus, as standard process for steady state CFD analysis. An overview of the computational domain is shown in Figure 1, showing both the geometry of the airfoils as well as the axial velocity contours at the mixing plane locations. The computational mesh was generated using Numeca AUTOGRID5, and the CFD solver used was Trace, with SST $K-\omega$ turbulence model. More details on the computational setup used have been provided in previous publications [10]. The development of build-specific CFD models for compressor aerodynamic predictions and its validation against engine test data was also presented in previous work by the authors, discussing the relative effects of various manufacturing and build variations [1]. The quantification of the accuracy of the CFD simulations in comparison to test data is outside of the scope of the current work. The aim of the proposed architecture is to provide predictions of the flow field and overall performance of the compressor with accuracy similar to that of the CFD model, but without the associated computational cost.

### v5

@@ Abstract
Applications of deep learning to physical simulations such as Computational Fluid Dynamics have recently experienced a surge in interest, and their viability has been demonstrated in different domains. However, due to the highly complex, turbulent and three-dimensional flows, they have not yet been proven usable for turbomachinery applications. Multi-stage axial compressors for gas turbine applications represent a remarkably challenging case, due to the high-dimensionality of the regression of the flow-field from geometrical and operational variables. This paper demonstrates the development and application of a deep learning framework for predictions of the flow field and aerodynamic performance of multi-stage axial compressors. A physics-based dimensionality reduction approach unlocks the potential for flow-field predictions, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. Compared to traditional "black-box" surrogate models, it provides explainability to the predictions of the overall performance by identifying the corresponding aerodynamic drivers. The model is applied to manufacturing and build variations, as the associated performance scatter is known to have a significant impact on $CO_{2}$ emissions, which poses a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark, in real-time, for an industrially relevant application. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance with actionable and explainable data.
@@
articletype: RESEARCH ARTICLEjournal: Pre-print submitted to Data-Centric Engineering
{Frontmatter}
@@
Bruni et al.
[…]
{policy}
[Impact Statement] This paper demonstrates the development and application of a deep learning framework for predictions of the flow-field and aerodynamic performance of multi-stage axial compressors. A physics-based dimensionality reduction unlocks the potential to use supervised learning approaches for turbomachinery flow field predictions, while providing explainability to the overall performance predictions. The model is applied to manufacturing and build variations, as the associated scatter in compressor efficiency is known to have a significant impact on the overall performance and $CO_{2}$ emissions of the gas turbine, thus posing a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance with actionable and explainable data, and reduce requirements for expensive physical tests.
@@ 1 Introduction
The highly complex, turbulent and three-dimensional flow-field of multi-stage axial compressors has traditionally been one the main challenges for the application of CFD (Computational Fluid Dynamics) in the turbomachinery industry. In this context, the flow field consists of the aerodynamic variables of interest (such as pressure, temperature, velocities, etc.) predicted on a numerical discretization of the domain (i.e. mesh or grid), solving the Reynolds-Averaged Navier Stokes equations (RANS). Advancements in modelling and numerical methods have progressively improved accuracy and computational cost, making CFD analyses an integral part of industrial design processes, greatly improving the design capabilities of gas turbine manufacturers. This has led to an interest in the development of robust design strategies to account for real world effects in the design and analyses, which are known to affect the overall performance as well as the operability of the compressor. Gas turbine manufacturers have gathered vast amount of operational data over the years, enabling them to establish best practices and design guidelines for manufacturing and build processes. These guidelines, often consider a trade-off between product cost and engine performance, with tolerance ranges specified accordingly. However, the impact on performance, used to define these ranges, is typically based on previous experience and simplified correlations. The identification of manufacturing and build uncertainties impact on the operational performance can be used to both drive improvements in the manufacturing technology, and inform the design strategies to produce robust designs that are desensitized to these variabilities.
1.1 The challenge
@@
Characterizing the impact of all the relevant manufacturing and build variations which are expected to affect the overall performance of a given compressor design, can be achieved by engine build-specific CFD models. While having demonstrated to provide excellent agreement with test data Bruni and Krishnababu (2024), these models are not traditionally used as part of the manufacturing and build process as a day-to-day occurrence, due to the associated computational cost and requirement for specialized engineers to carry out the analyses. Therefore, there is a technology gap for the development of predictive models that can match the accuracy of CFD, but with quicker turn-around times.
1.2 Our contribution
@@
This paper demonstrates the development and application of C(NN)FD, a deep learning framework previously introduced by the authors Bruni et al. (2024) for predictions of the flow field and aerodynamic performance of turbomachinery. The scalability is demonstrated from a proof of concept to a significantly more computationally expensive use-case of great industrial relevance, multi-stage axial compressors for gas turbine applications. We propose a physics-based dimensionality reduction to unlock the potential to use supervised learning approaches for flow-field predictions, as it re-formulates the regression problem from un-structured to structured, as well as reducing the number of degrees of freedom. The proposed architecture also has a competitive advantage compared to traditional "black-box" surrogate models, as it provides explainability to the predictions of variations in the overall performance. The corresponding aerodynamic drivers are identified by providing predictions for the 3D "structured" flow-field, as well as radial distributions at all the inter-row locations, 1D row-wise distributions and stage-wise distributions. The model is applied to manufacturing and build variations, as the associated scatter in efficiency is known to have a significant impact on the overall performance and $CO_{2}$ emissions of the gas turbine, therefore posing a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark, in real-time, for an industrially relevant application. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance with actionable and explainable data. This unlocks the capability to model each engine build avoiding the high computational costs, long lead times, specialized personnel requirements of traditional CFD analyses, and reducing the requirement for expensive physical tests. An overview of the C(NN)FD framework is provided in Figure 1.
Figure 1: Overview of the C(NN)FD framework
2 Literature Review
@@ 2.1 Effect of manufacturing and build variations on the aerodynamic performance
The impact of manufacturing and build variations on the overall performance of gas turbines is known to be significant Montomoli (2018) Wang and Zheng (2020), with the axial compressor being a large contributor. For a given design, compressor efficiency is affected by a combination of in-tolerance variations, which can lead to a substantial increase in $CO_{2}$ emissions compared to the design target. Therefore, accurately predicting the performance of each engine build by modelling the impact of manufacturing and build variations on engine performance within a short timescale is of great industrial and environmental importance. The focus of the current work is on tip clearance and surface roughness variations, two of the main sources of performance variability, but the framework is readily generalisable to other manufacturing variations. A detailed review of the effect of the various manufacturing and build variations on the compressor aerodynamic performance, and the associated modelling strategies has been previously been provided by the authors Bruni and Krishnababu (2024). In the following sections, a brief overview is included to provide the readers with the necessary background to interpret the results presented in this paper.
@@ 2.1.1 Tip Clearance variation background
The term tip clearance refers to the radial distance between the stationary and rotating components within the compressor, such as the rotor blades and casing, or stator blades and rotor hub. Generally, the tip clearance must be large enough to prevent rubbing during operation, but small enough to avoid excessive aerodynamic losses. Axial compressors are designed to operate at a defined tip clearance, which can vary across stages. During manufacturing, the "cold clearance" is specified and measured as part of the build process in stationary conditions. However, during operation, the "hot clearance" can be significantly smaller due to thermal expansion and centrifugal effects. The cold clearance values have a target value and a tolerance range, set by gas turbine manufacturers, based on trade-offs between undesirable rubbing and minimizing tip leakage losses. For a given "cold clearance", the corresponding "hot clearance" can be calculated using either analytical models or operational experience. The "hot clearance" is the actual clearance found during operation, and therefore is used in the CFD and machine learning models presented in this paper. The aerodynamic mechanisms driving tip leakage flows have been extensively researched over the years Storer and Cumpsty (1991) and can be modeled analytically with varying degrees of fidelity. Standard practice in both academia and industry, is to use steady-state single passage CFD calculations to model the effects of tip clearance variations on the overall performance Sakulkaew et al. (2013), which are used as ground-truth for this study. This assumes that circumferential variations in tip clearance can be neglected. Overall, non-axisymmetric tip clearance variations are known to affect the overall performance and surge margin Suriyanarayanan et al. (2021) and are of relevance especially for in-service degradation. However, for manufacturing variations, they can be neglected Bruni et al. (2024) and single passage models can be used for the analyses.
@@ 2.1.2 Surface roughness variation background
Traditionally, CFD analyses are performed assuming aerofoil and gas path surfaces to be aerodynamically smooth. However, this assumption often does not hold true for production components, which typically exhibit non-negligible roughness due to coating selection or manufacturing methods. A trade-off exists between production cost and surface roughness, as achieving smoother finishes tends to be progressively more expensive. Therefore, it is crucial to characterize the impact of different surface roughness levels on aerodynamic performance Bruni and Krishnababu (2024). This is particularly relevant when comparing engine test data from different builds with coatings from various suppliers, which can exhibit significantly different roughness levels. Experimental studies have demonstrated that increased roughness on the aerofoil surface adversely affects efficiency Suder et al. (1995). Additionally, surface roughness has been found to influence flow separation topology, leading to increased hub end-wall losses due to three-dimensional suction side separation Gbadebo et al. (2004). One popular method to model roughness analytically is to relate the centerline average roughness ($R_{a}$) to the equivalent sand grain roughness size ($k_{s}$). Various correlations have been proposed in the literature, with values ranging between $102<k_{s}/R_{a}<$ 10 Bons (2010). Specifying a constant roughness value has been shown to provide reasonable agreement with test data for manufactured blades, particularly in transonic compressors Kozulovic et al. (2022). A parametric study to identify the most relevant correlation for gas turbine high Reynolds number applications Hummel et al. (2005), resulted in the expression $k_{s}=5.2R_{a}$, which is utilised in this work.
@@ 2.2 Machine Learning applications to turbomachinery
In recent years, there has been a significant amount of research into the utilisation of machine learning techniques for turbomachinery applications. These include predicting the aerodynamic performance due to tip clearance variations Krishnababu et al. (2021) and design modifications Pongetti et al. (2021), as well as aeromechanic predictions for both forced response Bruni et al. (2022) and flutter He et al. (2022). As often the output variables are predicted directly from the input variables with a "black box" approach, these methodologies have varying degrees of accuracy and generalisability, depending on the complexity of the application. However, the ability to accurately predict variations in target variables for turbomachinery in a manner that is both reliable and generalisable, is contingent upon the accuracy of flow field predictions. It is only through precise flow field predictions that the overall performance figure of interest can be calculated through the application of relevant physical equations that govern the dynamics. The potential benefits of using machine learning approaches for predicting full flow fields, have been demonstrated in some simplified cases in the literature for 2D aerofoils on both cartesian Thuerey et al. (2020) Chen and Thuerey (2023) Ribeiro et al. (2020) and unstructured Kashefi et al. (2021) grids. CNNs have been demonstrated to predict the flow field downstream of a single fan row Li et al. (2023), which was then used for noise predictions applications. A similar approach was used also by Rao et. al. Rao et al. (2023) for the semantic segmentation of an aero-engine intake using U-Net and U-Net++ architectures. Graph neural networks Harsch and Riedelbauch (2021) Pfaff et al. (2021) have also been applied to 3D turbomachinery, with a single row compressor Beqiraj et al. (2022) and turbine Li et al. (2022), which suggested potential challenges related to scalability and practicality for industrial applications. For instance, the computational cost associated with training, and the size of the required model, make the approach impractical for multi-stage compressors. These applications typically use computational meshes in the order of magnitude of 10 million nodes. Furthermore, storing the CFD data required for training and future simulations would require several petabytes, for a typical engine manufacturer, which is not practical. Purely data-driven approaches typical of supervised learning require large amount of data - so that the network can learn the solution manifold. A recent review Thuerey et al. (2021) has highlighted the potential of other approaches such as Physics-Informed Neural Networks (PINNs), Reinforcement Learning (RL), and Generative Adversarial Networks (GANs) Chu et al. (2021) to mitigate these issues. However, none of those methods can currently provide accurate, generalisable and scalable solutions to complex 3D turbomachinery applications.
3 Novel contribution: Physics-based dimensionality reduction
@@
To address these challenges, we introduce a pre-processing step to extract only the relevant engineering data from the 3D computational domain, such as: blade-to-blade planes, blade surfaces, axial cuts, loading distributions, radial distributions, stage-wise performance, and overall performance. In most cases, these provide sufficient information to the aerodynamicist for design and analyses activities, while close examination of the 3D flow field of a CFD solution is generally only necessary in non-standard cases. Once the relevant engineering data has been extracted from the computational domain, it can be interpolated onto a simplified grid with a resolution that is deemed acceptable for the application in question. The regression problem is therefore reformulated, from predicting the relevant variables on an unstructured computational grid in the order of magnitude of $\sim 10^{7}$ nodes, to a structured grid in the order of magnitude of $\sim 10^{5}$ nodes. Only selected fundamental variables are predicted, while the derived ones can be computed using the appropriate physical equations, significantly reducing the computational cost without sacrificing accuracy, as only the necessary data required to calculate the target variables is retained. Targeting the flow field predictions and using the relevant physical equations to calculate the corresponding overall performance, renders the methodology generalisable, while filtering only relevant parts of the CFD solution makes the methodology scalable to complex industrial applications. This physics-based dimensionality reduction, unlocks the potential to use supervised learning approaches for turbomachinery applications, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. The authors first introduced this approach in a previous work for a single stage compressor application Bruni et al. (2024), which is now extended to a multi-stage compressor application, with a computational domain one order of magnitude larger. The framework developed in the previous paper is demonstrated to be translated from an academic test case, to a complex real-world application with significant engineering relevance. While the focus of the current work is on multi-stage axial compressors, the framework is directly applicable to any other turbomachinery application, such as axial turbines.
@@ 4.1 Data Generation
The ground-truth data used for training are the CFD results for the configuration of interest. This computational model consisted of 10 stages, with mixing plane interfaces. To reduce the computational cost of the CFD analysis, rotational periodicity is assumed, and only a single passage needs to be considered instead of the full annulus, as standard process for steady state CFD analysis. An overview of the computational domain is shown in Figure 2, showing both the geometry of the airfoils as well as the axial velocity contours at the mixing plane locations. The computational mesh was generated using Numeca AUTOGRID5, and the CFD solver used was Trace, with SST $K-\omega$ turbulence model. More details on the computational setup used have been provided in previous publications Bruni et al. (2022). The development of build-specific CFD models for compressor aerodynamic predictions and its validation against engine test data was also presented in previous work by the authors, discussing the relative effects of various manufacturing and build variations Bruni and Krishnababu (2024). The quantification of the accuracy of the CFD simulations in comparison to test data is outside of the scope of the current work. The aim of the proposed architecture is to provide predictions of the flow field and overall performance of the compressor with accuracy similar to that of the CFD model, but without the associated computational cost.